A train covers 60 miles between 2 p.m. and 4 p.m. How fast was it going at 3 p.m.?

Grade:11

2 Answers

Vikas TU
4 years ago
Dear student 
speed = distance/ time
here speed= 60/2= 30 miles per hr
30 miles per hr
It is the ans 
Hope tis helps 
Good Luck 
Arun
4 years ago
Dear student

The speed is traveled distance (60 miles) divided by traveled time (4pm – 2pm = 2hours):

60/2 = 30 miles per hour

 

 

Hope it helps
 
Regards
Arun